### v2

# Toric ideals of matching polytopes and edge colorings

Abstract.
In the present paper, we investigate the maximal degree of minimal generators of the toric ideal of the matching polytope of a graph. It is known that the toric ideal associated to a bipartite graph is generated by binomials of degree at most $3$. We show that this fact is equivalent to a result in the theory of edge colorings of bipartite multigraphs. Moreover, a characterization of bipartite graphs whose toric ideals are generated by quadratic binomials is given. Finally, we discuss the maximal degree of minimal generators of the toric ideal associated to a general graph and give a conjecture.
Key words and phrases:
matching polytope, Birkhoff polytope, flow polytope, stable set polytope, toric ideal, coloring
2020 Mathematics Subject Classification:
05C15, 13P10, 13F65
1. Introduction
In the present paper, we discuss a relationship between an algebraic property of toric ideals arising from graphs and a combinatorial property of edge colorings of multigraphs. Throughout this paper, we assume that a graph is simple, namely, it has no loops and no multiple edges and, a multigraph has no loops. Let $G$ be a graph with the vertex set $V(G)=[d]:=\{1,2,\ldots,d\}$ and the edge set $E(G)=\{e_{1},\dots,e_{n}\}$. A matching of $G$ is a set of pairwise non-adjacent edges of $G$, and a perfect matching of $G$ is a matching that covers every vertex of $G$. Let $M(G)$ (resp. $PM(G)$) denote the set of all matchings (resp. perfect matchings) of $G$. Given a subset $M\subset E(G)$, we associate the $(0,1)$-vector $\rho(M)=\sum_{e_{j}\in M}{\mathbf{e}}_{j}\in{\mathbb{R}}^{n}$. Here ${\mathbf{e}}_{j}$ is the $j$th unit coordinate vector in ${\mathbb{R}}^{n}$. For example, $\rho(\emptyset)=(0,\ldots,0)\in{\mathbb{R}}^{n}$. Then the (full) matching polytope $\mathcal{M}_{G}$ of $G$ is defined as the convex hull
$\mathcal{M}_{G}=\operatorname{conv}\left\{\rho(M):M\in M(G)\right\}$
and the perfect matching polytope $\mathcal{P}_{G}$ of $G$ is defined as
$\mathcal{P}_{G}=\operatorname{conv}\left\{\rho(M):M\in PM(G)\right\}.$
Note that $\mathcal{P}_{G}$ is a face of ${\mathcal{M}}_{G}$. Moreover, the perfect matching polytope of a complete bipartite graph $K_{d,d}$ is called the Birkhoff polytope, denoted by $\mathcal{B}_{d}$.
In [8], it was conjectured that the toric ideal $I_{\mathcal{B}_{n}}$ of the Birkhoff polytope $\mathcal{B}_{n}$ is generated by binomials of degree at most $3$, and this conjecture was shown in [22]. Moreover, in [9], by using this result, the toric ideal of a flow polytope is generated by binomials of degree at most $3$. For a homogeneous ideal $I$, let $\omega(I)$ denote the maximal degree of minimal generators of $I$. Since the matching polytope of a bipartite graph is unimodularly equivalent to a flow polytope (see Appendix A), the following result holds:
Theorem 1.1 ([9]).
For a bipartite graph $G$, one has $\omega(I_{{\mathcal{M}}_{G}})\leq 3$.
Next, we recall a result of edge-colorings of multigraphs. Let $G$ be a multigraph. For a $k$-edge-coloring $f$ of $G$ and a color $1\leq j\leq k$, let $M^{(e)}(f,j)$ denote the set of all edges of color $j$. We say that two $k$-edge-colorings $f$ and $g$ of $G$ differ by an $m$-colored subgraph if there is a set of colors $S$ of size $m$ such that $M^{(e)}(f,j)\neq M^{(e)}(g,j)$ for each $j\in S$, but $M^{(e)}(f,j)=M^{(e)}(g,j)$ for each $j\notin S$. For two $k$-edge-colorings $f,g$ of $G$, we write $f\sim_{r}g$ if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-edge-colorings of $G$ with $f_{0}=f$ and $f_{s}=g$ such that $f_{i}$ differs from $f_{i-1}$ by a $k_{i}$-colored subgraph with $k_{i}\leq r$. Note that $f\sim_{r}g$ implies $f\sim_{r+1}g$. In [1, 2, 3], the following result was shown:
Theorem 1.2 ([1, 2, 3]).
Let $G$ be a bipartite multigraph. Then for any $k$-edge-colorings $f$ and $g$ of $G$, one has $f\sim_{3}g$.
In the present paper, we show that Theorems 1.1 and 1.2 are equivalent. For a simple graph $G$ on $[d]$ with $E(G)=\{e_{1},e_{2},\ldots,e_{n}\}$ and ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}^{n}_{\geq 0}$, let $G^{(e)}_{{\mathbf{a}}}$ be the multigraph on $[d]$ such that $G^{(e)}_{{\mathbf{a}}}$ has $a_{i}$ multiedges $e_{i}$ for each $i$. We call $G^{(e)}_{{\mathbf{a}}}$ the edge-replication multigraph of $G$ on $\mathbf{a}$}. Then our main result is the following:
Theorem 1.3.
Let $G$ be a graph with $n$ edges. Then $\omega(I_{{\mathcal{M}}_{G}})\leq r$ if and only if for any ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$ and for any $k$-edge-colorings $f$ and $g$ of $G^{(e)}_{{\mathbf{a}}}$, one has $f\sim_{r}g$.
Since any edge-replication multigraph of a simple bipartite graph is bipartite, Theorems 1.1 and 1.2 are equivalent from this theorem. In particular, this gives an alternative proof of the conjecture of [8] mentioned above. Moreover, we can show a similar result for perfect matching polytopes (Proposition 6.8).
On the other hand, we give a characterization of a bipartite graph such that $\omega(I_{{\mathcal{M}}_{G}})=2$, i.e., $I_{{\mathcal{M}}_{G}}$ is generated by quadratic binomials. In fact,
Theorem 1.4.
Let $G$ be a bipartite graph. Then the following conditions are equivalent:
  • (i)
    $\omega(I_{{\mathcal{M}}_{G}})=2$;
  • (ii)
    $G$ has no odd subdivision of $K_{2,3}$ as a subgraph;
  • (iii)
    each block of $G$ is a bipartite graph having no odd subdivision of $K_{2,3}$ as a subgraph.
Otherwise, one has $\omega(I_{{\mathcal{M}}_{G}})=3$.
Finally, we discuss the maximal degree of minimal generators of $I_{\mathcal{M}_{G}}$ for a general graph $G$. There exists a non-bipartite graph $G$ with $\omega(I_{\mathcal{M}_{G}})=4$ (see Example 6.1). Hence we cannot extend Theorem 1.2 to the case of arbitrary graphs. However, motivated by a conjecture in [1], we give the following conjectures:
Conjecture 1.5.
Let $G$ be a (non-bipartite) graph. Then one has $\omega(I_{{\mathcal{M}}_{G}})\leq 4$.
Conjecture 1.6.
Let $G$ be a (non-bipartite) graph. Then one has $\omega(I_{{\mathcal{P}}_{G}})\leq 4$.
In general, we have $\omega(I_{{\mathcal{P}}_{G}})\leq\omega(I_{{\mathcal{M}}_{G}})$ (Proposition 6.9 (a)). On the other hand, there exists a graph $G$ such that $\omega(I_{{\mathcal{M}}_{G}})\neq\omega(I_{{\mathcal{P}}_{G}})$. However, we can show that these conjectures are equivalent (Proposition 6.10).
The present paper is organized as follows: In Section 2, we introduce the toric ideal of a lattice polytope and define the stable set polytope associated with a graph. Note that ${\mathcal{M}}_{G}$ is the stable set polytope of the line graph of $G$. Section 3 proves Theorem 1.3 by providing an upper bound on the maximal degree of minimal generators of the toric ideal of a stable set polytope using vertex-colorings. Section 4 gives an algebraic method to determine if $f\sim_{r}g$ for $k$-vertex-colorings $f$ and $g$ of a graph $G$. In Section 5, we extend Theorem 1.1 to the case of line perfect graphs and prove Theorem 1.4 by characterizing line perfect graphs such that the toric ideals of their matching polytopes are generated by quadratic binomials. Finally, Section 6 considers examples of $I_{{\mathcal{M}}_{G}}$ for general graphs and Conjecture 1.5 for graphs with a small number of vertices, and discusses perfect matching polytopes.
2. Preliminaries
In this section, we introduce the toric ideal of a lattice polytope and define the stable set polytope associated with a graph.
2.1. Toric rings and toric ideals
Let ${\mathcal{P}}\subset{\mathbb{R}}^{d}_{\geq 0}$ be a lattice polytope with ${\mathcal{P}}\cap{\mathbb{Z}}^{d}_{\geq 0}=\{{\mathbf{a}}_{1},\ldots,{\mathbf{a}}_{n}\}$ and let ${\mathbb{K}}[{\mathbf{t}},s]:={\mathbb{K}}[t_{1},\ldots,t_{d},s]$ be the polynomial ring in $d+1$ variables over a field $\mathbb{K}$}. Given a nonnegative integer vector ${\mathbf{a}}=(a_{1},\ldots,a_{d})\in{\mathbb{Z}}_{\geq 0}^{d}$, we write ${\mathbf{t}}^{{\mathbf{a}}}:=t_{1}^{a_{1}}t_{2}^{a_{2}}\cdots t_{d}^{a_{d}}\in{\mathbb{K}}[{\mathbf{t}},s]$. The toric ring of $\mathcal{P}$} is
${\mathbb{K}}[{\mathcal{P}}]:={\mathbb{K}}[{\mathbf{t}}^{{\mathbf{a}}_{1}}s,\ldots,{\mathbf{t}}^{{\mathbf{a}}_{n}}s]\subset{\mathbb{K}}[{\mathbf{t}},s].$
We regard ${\mathbb{K}}[{\mathcal{P}}]$ as a homogeneous algebra by setting each $\deg({\mathbf{t}}^{{\mathbf{a}}_{i}}s)=1$. Let $R[{\mathcal{P}}]={\mathbb{K}}[x_{1},\ldots,x_{n}]$ denote the polynomial ring in $n$ variables over $\mathbb{K}$} with each $\deg(x_{i})=1$. The toric ideal $I_{{\mathcal{P}}}$ of $\mathcal{P}$} is the kernel of the surjective homomorphism $\pi:R[{\mathcal{P}}]\to{\mathbb{K}}[{\mathcal{P}}]$ defined by $\pi(x_{i})={\mathbf{t}}^{{\mathbf{a}}_{i}}s$ for $1\leq i\leq n$. Note that $I_{{\mathcal{P}}}$ is a prime ideal generated by homogeneous binomials. The toric ring ${\mathbb{K}}[{\mathcal{P}}]$ is called quadratic if $I_{\mathcal{P}}$ is generated by quadratic binomials. We say that $I_{\mathcal{P}}$ is generated by quadratic binomials” even if $I_{\mathcal{P}}=\{0\}$. In particular, $\omega(I_{\mathcal{P}})\geq 2$ and $\omega(\{0\})=2$. The following is known.
Proposition 2.1 ([16], [15, Theorem 1.3]).
Let $F$ be a face of a lattice polytope $\mathcal{P}$}. If $\mathcal{G}$} is a set of generators of $I_{\mathcal{P}}$, then ${\mathcal{G}}\cap R[F]$ is a set of generators of $I_{F}$. In particular, we have $\omega(I_{F})\leq\omega(I_{\mathcal{P}})$.
2.2. Stable set polytopes and stable set ideals
Let $G$ be a simple graph on $[d]$. A subset $S\subset[d]$ is called a stable set (or an independent set) of $G$ if $\{i,j\}\notin E(G)$ for all $i,j\in S$ with $i\neq j$. In particular, the empty set $\emptyset$ and any singleton $\{i\}$ with $i\in[d]$ are stable. Let $S(G)$ denote the set of all stable sets of $G$. Then the stable set polytope ${\mathcal{S}}_{G}$ of $G$ is defined as
${\mathcal{S}}_{G}=\operatorname{conv}\{\rho(S):S\in S(G)\}.$
The matching polytope ${\mathcal{M}}_{G}$ is a stable set polytope of some graph. Indeed, the line graph $L(G)$ of $G$ is a simple graph whose vertex set is $E(G)$ and whose edge set is
[…]
Then one has ${\mathcal{M}}_{G}={\mathcal{S}}_{L(G)}$ by changing coordinates.
For a graph $G$, the toric ideal $I_{{\mathcal{S}}_{G}}$ is called the stable set ideal of $G$. We can describe a system of generators of $I_{{\mathcal{S}}_{G}}$ in terms of $k$-vertex-colorings. Given a graph $G$ on the vertex set $[d]$, and ${\mathbf{a}}=(a_{1},\ldots,a_{d})\in{\mathbb{Z}}_{\geq 0}^{d}$, let $G_{\mathbf{a}}$ be the graph obtained from $G$ by replacing each vertex $i\in[d]$ with a complete graph $G^{(i)}$ of $a_{i}$ vertices (if $a_{i}=0$, then just delete the vertex $i$), and joining all vertices $x\in G^{(i)}$ and $y\in G^{(j)}$ such that $\{i,j\}$ is an edge of $G$. In particular, if ${\mathbf{a}}=(1,\ldots,1)$, then $G_{\mathbf{a}}=G$. If ${\mathbf{a}}={\bf 0}$, then $G_{\mathbf{a}}$ is the null graph (a graph without vertices). In addition, if $\mathbf{a}$} is a $(0,1)$-vector, namely, ${\mathbf{a}}\in\{0,1\}^{d}$, then $G_{\mathbf{a}}$ is an induced subgraph of $G$. We call $G_{\mathbf{a}}$ a vertex-replication graph of $G$. Given a $k$-vertex-coloring $f$ of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$, we associate $f$ with a monomial
${\mathbf{x}}_{f}:=x_{S_{i_{1}}}\cdots x_{S_{i_{k}}}\in R[{\mathcal{S}}_{G}],$
where $S_{i_{\ell}}=\{j\in[d]:G^{(j)}\cap f^{-1}(\ell)\neq\emptyset\}$ for $\ell=1,2,\ldots,k$. Conversely, let $m=x_{S_{i_{1}}}\cdots x_{S_{i_{k}}}\in R[{\mathcal{S}}_{G}]$ be a monomial of degree $k$. Then, for ${\mathbf{a}}=(a_{1},\ldots,a_{n})$ with $a_{p}=|\{\ell:p\in S_{i_{\ell}}\}|$, there exists a $k$-vertex-coloring $f$ of $G_{{\mathbf{a}}}$ such that ${\mathbf{x}}_{f}=m$ (see [18, Lemma 3.2]). Note that, for $k$-vertex-colorings $f$ and $g$ of an induced subgraph of $G$, ${\mathbf{x}}_{f}={\mathbf{x}}_{g}$ if and only if $g$ is obtained from $f$ by permuting colors. In this paper, we identify $f$ and $g$ if $g$ is obtained from $f$ by permuting colors. Then we can describe a system of generators of $I_{{\mathcal{S}}_{G}}$ as follows:
Proposition 2.2 ([18, Theorem 3.3]).
Let $G$ be a simple graph on $[d]$. Then one has
[…]
where $\chi(G_{{\mathbf{a}}})$ is the chromatic number of $G_{{\mathbf{a}}}$.
If $G^{\prime}$ is an induced subgraph of a graph $G$, then ${\mathcal{S}}_{G^{\prime}}$ is a face of ${\mathcal{S}}_{G}$. From Proposition 2.1, we have the following.
Proposition 2.3.
Let $G^{\prime}$ be an induced subgraph of $G$. Then we have $\omega(I_{{\mathcal{S}}_{G^{\prime}}})\leq\omega(I_{{\mathcal{S}}_{G}})$.
If $G^{\prime}$ is a subgraph of $G$, then $L(G^{\prime})$ is an induced subgraph of $L(G)$. Hence the following fact holds from ${\mathcal{M}}_{G}={\mathcal{S}}_{L(G)}$.
Proposition 2.4.
Let $G^{\prime}$ be a subgraph of $G$. Then we have $\omega(I_{{\mathcal{M}}_{G^{\prime}}})\leq\omega(I_{{\mathcal{M}}_{G}})$.
3. A Bound on $\omega(I_{{\mathcal{S}}_{G}})$
In this section, we prove Theorem 1.3 by providing an upper bound on the maximal degree of minimal generators of the toric ideal of a stable set polytope using vertex-colorings.
Let $G$ be a graph. For a $k$-vertex-coloring of $G$ and a color $1\leq j\leq k$, let $M(f,j)$ denote the set of all vertices of color $j$. We say that two $k$-vertex-colorings $f$ and $g$ of $G$ differ by an $m$-colored subgraph if there is a set of colors $S$ of size $m$ such that $M(f,j)\neq M(g,j)$ for each $j\in S$, but $M(f,j)=M(g,j)$ for each $j\notin S$. For two $k$-vertex-colorings $f,g$ of $G$, we write $f\sim_{r}g$ if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-vertex-colorings of $G$ with $f_{0}=f$ and $f_{s}=g$ such that $f_{i}$ differs from $f_{i-1}$ by a $k_{i}$-colored subgraph with $k_{i}\leq r$. We give a proof of Theorem 1.3 by showing the following more general result.
Theorem 3.1.
Let $G$ be a graph on $[d]$. Then $\omega(I_{{\mathcal{S}}_{G}})\leq r$ if and only if for any ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ and for any $k$-vertex-colorings $f$ and $g$ of $G_{{\mathbf{a}}}$, one has $f\sim_{r}g$.
Proof.
(only if) Let $f$ and $g$ be $k$-vertex-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ and $k>r$. Then the binomial $F:={\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ belongs to $I_{{\mathcal{S}}_{G}}$. If $F=0$, then $g$ is obtained from $f$ by permuting colors. Assume that $F\neq 0$. By the hypothesis, $F$ is generated by binomials of degree at most $r$ in $I_{{\mathcal{S}}_{G}}$. From the theory of binomial ideals [11, Lemma 3.8], there exists an expression
(3.1) $F=\sum_{i=1}^{s}{\bf x}^{{\mathbf{w}}_{i}}({\mathbf{x}}_{f_{i}}-{\mathbf{x}}_{g_{i}}),$
where for each $i$, $f_{i}$ and $g_{i}$ are $k_{i}$-vertex-colorings of $G_{{\mathbf{a}}_{i}}$ with ${\mathbf{a}}_{i}\in{\mathbb{Z}}_{\geq 0}^{d},k_{i}\leq r$ and ${\mathbf{x}}_{f_{i}}-{\mathbf{x}}_{g_{i}}$ $(\neq 0)$ is an irreducible binomial. We may suppose that ${\mathbf{x}}_{f}={\mathbf{x}}^{{\mathbf{w}}_{1}}{\mathbf{x}}_{f_{1}}$ and ${\mathbf{x}}^{{\mathbf{w}}_{s}}{\mathbf{x}}_{g_{s}}={\mathbf{x}}_{g}$. Set ${\mathbf{x}}_{f_{1}}=x_{S_{1}}x_{S_{2}}\cdots x_{S_{k_{1}}},{\mathbf{x}}_{g_{1}}=x_{S^{\prime}_{1}}x_{S^{\prime}_{2}}\cdots x_{S^{\prime}_{k_{1}}}$ and ${\mathbf{x}}^{{\mathbf{w}}_{1}}=x_{T_{1}}x_{T_{2}}\cdots x_{T_{k-k_{1}}}$ with $S_{i},S^{\prime}_{i},T_{j}\in S(G)$. Since ${\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}}\in I_{{\mathcal{S}}_{G}}$, one has $S:=\bigcup_{1\leq i\leq k_{1}}S_{i}=\bigcup_{1\leq i\leq k_{1}}S^{\prime}_{i}$ as multisets. Moreover, it follows from ${\mathbf{x}}_{f}-{\mathbf{x}}^{{\mathbf{w}}_{1}}{\mathbf{x}}_{g_{1}}\in I_{{\mathcal{S}}_{G}}$ that there exists a $k$-vertex-coloring $g^{\prime}_{1}$ of $G_{{\mathbf{a}}}$ such that ${\mathbf{x}}_{g^{\prime}_{1}}={\mathbf{x}}^{{\mathbf{w}}_{1}}{\mathbf{x}}_{g_{1}}$. Then by exchanging colors and exchanging the coloring of vertices in each clique $G^{(j)}$ of $G_{{\mathbf{a}}}$ if necessary, we can assume that $M(f,j)\neq M(g^{\prime}_{1},j)$ for each $j\in S$ and $M(f,j)=M(g^{\prime}_{1},j)$ for each $j\notin S$. This implies that $g^{\prime}_{1}$ differs from $f$ by a $k_{1}$-colored subgraph. By performing this process repeatedly, we can obtain a sequence $g_{0}^{\prime},g_{1}^{\prime},\ldots,g_{s}^{\prime}$ of $k$-vertex-colorings of $G_{{\mathbf{a}}}$ with $g_{0}^{\prime}=f$ and ${\mathbf{x}}_{g_{s}^{\prime}}={\mathbf{x}}_{g}$ such that $g_{i}^{\prime}$ differs from $g_{i-1}^{\prime}$ by a $k_{i}$-colored subgraph with $k_{i}\leq r$. Then $g$ is obtained from $g_{s}^{\prime}$ by permuting colors. Hence one has $f\sim_{r}g$.
(if) Let $F={\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in I_{{\mathcal{S}}_{G}}$ where $f$ and $g$ are $k$-vertex-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ and $k>r$. From the assumption, there exists a sequence $f_{0},f_{1},\ldots,f_{t}$ of $k$-vertex-colorings of $G_{{\mathbf{a}}}$ with $f_{0}=f$ and $f_{t}=g$ such that $f_{i}$ differs from $f_{i-1}$ by a $k_{i}$-colored subgraph with $k_{i}\leq r$. Then there exists a set $S_{i}$ of colors with $|S_{i}|=k_{i}$ such that $M(f_{i},j)\neq M(f_{i-1},j)$ for each $j\in S_{i}$ and $M(f_{i},j)=M(f_{i-1},j)$ for each $j\notin S_{i}$. Let $f_{i}|_{S_{i}}$ be the $k_{i}$-vertex-coloring of the induced subgraph of $G_{{\mathbf{a}}}$ on the vertex set $f_{i}^{-1}(S_{i})$ induced from $f_{i}$. Then one has ${\mathbf{x}}_{f_{i-1}|_{S_{i}}}-{\mathbf{x}}_{f_{i}|_{S_{i}}}\in I_{{\mathcal{S}}_{G}}$. Similarly, let $f_{i}|_{\overline{S_{i}}}$ be the $(k-k_{i})$-vertex-coloring of the induced subgraph of $G_{{\mathbf{a}}}$ on the vertex set $\{j\in V(G_{\mathbf{a}}):f(j)\notin S_{i}\}$ induced from $f_{i}$. Then we obtain ${\mathbf{x}}_{f_{i-1}|_{\overline{S_{i}}}}={\mathbf{x}}_{f_{i}|_{\overline{S_{i}}}},{\mathbf{x}}_{f_{i-1}|_{\overline{S_{i}}}}{\mathbf{x}}_{f_{i-1}|_{S_{i}}}={\mathbf{x}}_{f_{i-1}}$ and ${\mathbf{x}}_{f_{i}|_{\overline{S_{i}}}}{\mathbf{x}}_{f_{i}|_{S_{i}}}={\mathbf{x}}_{f_{i}}$. Hence one has
$F={\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ $=({\mathbf{x}}_{f_{0}}-{\mathbf{x}}_{f_{1}})+({\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{f_{2}})+\cdots+({\mathbf{x}}_{f_{t-1}}-{\mathbf{x}}_{f_{t}})$
$={\mathbf{x}}_{f_{1}|_{\overline{S_{1}}}}({\mathbf{x}}_{f_{0}|_{S_{1}}}-{\mathbf{x}}_{f_{1}|_{S_{1}}})+{\mathbf{x}}_{f_{2}|_{\overline{S_{2}}}}({\mathbf{x}}_{f_{1}|_{S_{2}}}-{\mathbf{x}}_{f_{2}|_{S_{2}}})+\cdots+{\mathbf{x}}_{f_{t}|_{\overline{S_{t}}}}({\mathbf{x}}_{f_{t-1}|_{S_{t}}}-{\mathbf{x}}_{f_{t}|_{S_{t}}}).$
Since ${\mathbf{x}}_{f_{i-1}|_{S_{i}}}-{\mathbf{x}}_{f_{i}|_{S_{i}}}\in I_{{\mathcal{S}}_{G}}$ is a binomial of degree $k_{i}\leq r$, $F$ is generated by binomials of degree $\leq r$. This implies $\omega(I_{{\mathcal{S}}_{G}})\leq r$. ∎
Proof of Theorem 1.3.
Let $G$ be a simple graph with $d$ edges and take ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$. Then the edge-replication multigraph $G^{(e)}_{{\mathbf{a}}}$ coincides with the vertex-replication graph $L(G)_{{\mathbf{a}}}$. Moreover, a $k$-edge-coloring $f$ of $G^{(e)}_{{\mathbf{a}}}$ can be regard as a $k$-vertex-coloring of $L(G)_{{\mathbf{a}}}$. In particular, two $k$-edge-colorings $f$ and $g$ of $G^{(e)}_{{\mathbf{a}}}$ differ by an $m$-colored subgraph if and only if $f$ and $g$ differ by an $m$-colored subgraph as $k$-vertex-colorings of $L(G)_{{\mathbf{a}}}$. Hence Theorem 1.3 follows from Theorem 3.1. ∎
4. $r$-coloring ideals
In this section, we give an algebraic method to determine if $f\sim_{r}g$ for two $k$-vertex-colorings $f,g$ of a graph $G$.
Given a $k$-vertex-coloring $f$ of $G$, and integers $1\leq i<j\leq k$, let $H$ be a connected component of the induced subgraph of $G$ on the vertex set $f^{-1}(i)\cup f^{-1}(j)$. Then we can obtain a new $k$-vertex-coloring $g$ of $G$ by setting
[…]
We say that $g$ is obtained from $f$ by a Kempe switching. Two $k$-vertex-colorings $f$ and $g$ of $G$ are called Kempe equivalent, if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-vertex-colorings of $G$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. It is easy to see that $f$ and $g$ are Kempe equivalent if and only if $f\sim_{2}g$. In [19], the third and fourth author introduced $2$-coloring ideals associated with graphs to examine when $f\sim_{2}g$. Given a graph $G$ on $[d]$, the $2$-coloring ideal is defined as follows:
[…]
Proposition 4.1 ([19, Theorem 1.1]).
Let $G$ be a graph and take two $k$-vertex-colorings $f,g$ of $G$. Then $f\sim_{2}g$ if and only if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G,2}$.
We generalize this result to determine $f\sim_{r}g$. Given an integer $r\geq 2$, we define the $r$-coloring ideal of $G$ as follows:
[…]
Note that $J_{G,r}\subset J_{G,r+1}$. For $r\geq|V(G)|$, one has $J_{G,r}=J_{G,r+1}$.
Theorem 4.2.
Let $G$ be a graph and take two $k$-vertex-colorings $f,g$ of $G$. Then $f\sim_{r}g$ if and only if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G,r}$.
Proof.
This follows by a similar argument as in the proof of Theorem 3.1. ∎
5. A bound on $\omega(I_{{\mathcal{M}}_{G}})$ for a line perfect graph
In this section, we give a proof of Theorem 1.4 by showing a more general result for line perfect graphs.
Let $G$ be a graph on $[d]$ with edge set $E(G)$. A subset $C\subset[d]$ is called clique of $G$ if for any $i,j\in C$ with $i\neq j$, $\{i,j\}\in E(G)$. Let $\omega(G)$ denote the maximum cardinality of cliques of $G$. A graph $G$ is called perfect if every induced subgraph $H$ of $G$ satisfies $\chi(H)=\omega(H)$. Perfect graphs were introduced by Berge in [4]. A hole is an induced cycle of length $\geq 5$ and an antihole is the complement of a hole. In [7], Chudnovsky, Robertson, Seymour and Thomas showed that a graph is perfect if and only if it has no odd holes and no odd antiholes. This result is called the strong perfect graph theorem.
A line perfect graph is a graph whose line graph is perfect. Note that every bipartite graph is line perfect. A characterization of line perfect graphs is known. A vertex $v$ of a connected graph $G$ is called a cut vertex if the graph obtained by the removal of $v$ from $G$ is disconnected. Given a graph $G$, a block of $G$ is a maximal connected subgraph of $G$ with no cut vertices.
Proposition 5.1 ([12, 21]).
Let $G$ be a graph. Then the following conditions are equivalent:
  • (i)
    $G$ is line perfect;
  • (ii)
    $G$ has no odd cycle of length $\geq 5$ as a subgraph;
  • (iii)
    each block of $G$ is either a bipartite graph, $K_{4}$, or $K_{1,1,n}$.
The graph obtained by gluing two graphs at a clique $C$ of them is called a $|C|$-clique sum of them. (Here we do not remove any edges of the clique.) Clique sums of more than two graphs are defined by repeated application of this operation.
Proposition 5.2 ([13, Proposition 1]).
Suppose that $G$ is a clique sum of graphs $G_{1}$ and $G_{2}$. Then one has
$\omega(I_{{\mathcal{S}}_{G}})=\max\{\omega(I_{{\mathcal{S}}_{G_{1}}}),\omega(I_{{\mathcal{S}}_{G_{2}}})\}.$
Lemma 5.3.
Let $G$ be a graph whose blocks are $H_{1},\dots,H_{s}$. Then one has
$\omega(I_{{\mathcal{M}}_{G}})=\max\{\omega(I_{{\mathcal{M}}_{H_{1}}}),\dots,\omega(I_{{\mathcal{M}}_{H_{s}}})\}.$
Proof.
Since $G$ is a $1$-clique sum of $H_{1},\dots,H_{s}$, it is enough to show that
$\omega(I_{{\mathcal{M}}_{G}})=\max\{\omega(I_{{\mathcal{M}}_{G_{1}}}),\omega(I_{{\mathcal{M}}_{G_{2}}})\}$
if $G$ is a $1$-clique sum of $G_{1}$ and $G_{2}$ at a vertex $v$. Let $E_{i}=\{e\in E(G_{i}):v\in e\}$ for $i=1,2$. It then follows that $L(G)$ is a clique sum of $L(G_{1}),L(G_{2})$ and the complete graph $L(G^{\prime})$ along cliques $E_{1}$ and $E_{2}$, where $G^{\prime}$ is the graph whose edge set is $E_{1}\cup E_{2}$. Note that $I_{{\mathcal{M}}_{G^{\prime}}}=I_{{\mathcal{S}}_{L(G^{\prime})}}=\{0\}$. Then
$\omega(I_{{\mathcal{M}}_{G}})=\max\{\omega(I_{{\mathcal{M}}_{G_{1}}}),\omega(I_{{\mathcal{M}}_{G_{2}}}),\omega(I_{{\mathcal{M}}_{G^{\prime}}})\}=\max\{\omega(I_{{\mathcal{M}}_{G_{1}}}),\omega(I_{{\mathcal{M}}_{G_{2}}})\}$
from Proposition 5.2. ∎
As a generalization of Theorem 1.3 we give a bound of $\omega(I_{{\mathcal{M}}_{G}})$ for a line perfect graph $G$. In fact,
Theorem 5.4.
Let $G$ be a line perfect graph. Then one has $\omega(I_{{\mathcal{M}}_{G}})\leq 3$.
Proof.
From Proposition 5.1 and Lemma 5.3, it is enough to show that $\omega(I_{{\mathcal{M}}_{G}})\leq 3$ if $G$ is a bipartite graph, $K_{4}$, or $K_{1,1,n}$.
If $G$ is a bipartite graph, then $\omega(I_{{\mathcal{M}}_{G}})\leq 3$ by Theorem 1.1. If $G=K_{1,1,n}$, then $G$ is obtained from $K_{2,n}$ by adding a new edge $e$. Then $\{e\}$ is a unique matching of $K_{1,1,n}$ that contains $e$. Hence $I_{{\mathcal{M}}_{G}}$ and $I_{{\mathcal{M}}_{K_{2,n}}}$ have the same set of generators. Thus we have $\omega(I_{{\mathcal{M}}_{G}})=\omega(I_{{\mathcal{M}}_{K_{2,n}}})\leq 3$. Let $G=K_{4}$. Then $L(G)$ has no induced path with 4 vertices. Hence it is trivial that $L(G)$ is perfectly orderable, that is, there exists a linear order << on $V(L(G))$ such that no induced path with vertices $a,b,c,d$ and edges $\{a,b\},\{b,c\},\{c,d\}$ satisfies $a<b$ and $d<c$. It is known [17, Theorem 3.1] that $\omega(I_{{\mathcal{S}}_{G^{\prime}}})=2$ if $G^{\prime}$ is perfectly orderable. Thus $\omega(I_{{\mathcal{M}}_{G}})=\omega(I_{{\mathcal{S}}_{L(G)}})=2$. ∎
Combining this theorem and Theorem 1.3 we can obtain the following corollary.
Corollary 5.5.
Let $G$ be a multigraph whose underlying simple graph is line perfect. Then for any $k$-edge-colorings $f$ and $g$ of $G$, one has $f\sim_{3}g$.
Next, we characterize when $\omega(I_{{\mathcal{M}}_{G}})=2$ for a line perfect graph. Bertschi introduced a hereditary class of perfect graphs in [5]. An even pair in a graph $G$ is a pair of non-adjacent vertices of $G$ such that the length of all induced paths between them is even. Contracting a pair of vertices $\{x,y\}$ in a graph $G$ means removing $x$ and $y$ and adding a new vertex $z$ with edges to every neighborhood of $x$ or $y$. A graph $G$ is called even-contractile if there exists a sequence $G_{0},\ldots,G_{k}$ of graphs satisfying the following:
  1. (i)
    $G=G_{0}$;
  2. (ii)
    each $G_{i}$ is obtained from $G_{i-1}$ by contracting an even pair of $G_{i-1}$;
  3. (iii)
    $G_{k}$ is a complete graph.
A graph $G$ is called perfectly contractile if every induced subgraph of $G$ is even-contractile. Every perfectly contractile graph is perfect. In contrast to the strong perfect graph theorem, a forbidden graph characterization of perfectly contractile graphs is still open. However, there is a conjecture of this problem. An odd prism is a graph consisting of two disjoint triangles with three disjoint induced paths of odd length between them. Everett and Reed conjectured that a graph $G$ is perfectly contractile if and only if $G$ contains no odd holes, no antiholes and no odd prisms as induced subgraphs. On the other hand, the third and fourth authors and Shibata gave the following conjecture.
Conjecture 5.6 ([17, Conjecture 0.2]).
Let $G$ be a perfect graph. Then the following conditions are equivalent:
  • (i)
    $G$ is perfectly contractile;
  • (ii)
    $G$ contains no odd holes, no antiholes and no odd prisms;
  • (iii)
    $\omega(I_{{\mathcal{S}}_{G}})=2$.
A graph $G$ is called line perfectly contractile if its line graph $L(G)$ is perfectly contractile. If Conjecture 5.6 is true for a line perfect graph $G$, $\omega(I_{{\mathcal{M}}_{G}})=2$ if and only if $G$ is line perfectly contractile. We show that this claim is true by proving the following theorem which implies Theorem 1.4. An odd subdivision of a graph $G$ is a graph obtained by replacing each edge of $G$ by a path of odd length. Note that $G$ itself is an odd subdivision of $G$.
Theorem 5.7.
Let $G$ be a line perfect graph. Then the following conditions are equivalent:
  • (i)
    $\omega(I_{{\mathcal{M}}_{G}})=2$;
  • (ii)
    $G$ is line perfectly contractile;
  • (iii)
    $L(G)$ has no odd prisms;
  • (iv)
    $G$ has no odd subdivision of $K_{2,3}$ as a subgraph;
  • (v)
    each block of $G$ is either a bipartite graph having no odd subdivision of $K_{2,3}$ as a subgraph, $K_{3}$, $K_{4}$, or $K_{1,1,2}$.
Otherwise, $\omega(I_{{\mathcal{M}}_{G}})=3$.
In order to prove this theorem, we recall the following results.
Proposition 5.8 ([20]).
A dart-free graph is perfectly contractile if and only if it contains no odd holes, no antiholes and no odd prisms as induced subgraphs.
Proposition 5.9 ([18, Theorem 1.5 (a)]).
Let $G$ be a dart-free graph with no odd holes, no antiholes, and no odd prisms. Then one has $\omega(I_{{\mathcal{S}}_{G}})=2$.
Now, we give a proof of Theorem 5.7.
Proof of Theorem 5.7.
From [17, Theorem 1.7], (i) $\Rightarrow$ (iii) holds for any perfect graph.
Let $G$ be a line perfect graph. Since $L(G)$ is perfect, it has no odd holes and odd antiholes. In general, the line graph of a graph is dart-free (since claw-free) and has no graph $H$ below as an induced subgraph.
Since the complement of $H$ is the disjoint union of an edge and a path with 3 vertices, it follows that $L(G)$ has no antiholes of length $\geq 7$. Note that an antihole of length $6$ is an odd prism. Hence from Proposition 5.8, we have (ii) $\Leftrightarrow$ (iii). Moreover, from Proposition 5.9, we have (iii) $\Rightarrow$ (i). It is known that the line graph of an odd subdivision of $K_{2,3}$ is an odd prism. (A subdivision of $K_{2,3}$ is called theta.) Thus we have (iii) $\Leftrightarrow$ (iv). Finally, (iv) $\Leftrightarrow$ (v) follows from Proposition 5.1. ∎
Example 5.10.
Let $G$ be an outerplanar bipartite graph. It is known that $G$ has no $K_{2,3}$ as a minor. Hence the toric ring of the matching polytope of $G$ is quadratic.
6. A bound on $\omega(I_{{\mathcal{M}}_{G}})$ for a general graph
In this section, we consider $I_{{\mathcal{M}}_{G}}$ for a general graph $G$. First, we see examples of graphs $G$ with $\omega(I_{{\mathcal{M}}_{G}})=4$ by using Macaulay2 [10].
Example 6.1.
(1) Let $G_{1}$ be the graph as follows:
$G_{1}$
Using Macaulay2, one has $\omega(I_{{\mathcal{M}}_{G_{1}}})=4$. It then follows from Theorem 1.3 that there exist ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{8}$ and $4$-edge-colorings $f$ and $g$ of $(G_{1})^{(e)}_{{\mathbf{a}}}$ such that $f\not\sim_{3}g$. In fact, for the following two $4$-edge-colorings $f$ and $g$ of $G_{1}$, one has $f\not\sim_{3}g$.
$f$
$g$
Indeed, for any three (resp. two) colors, the subgraph consisting of all edges with the colors has a unique $3$-edge-coloring (resp. $2$-edge-coloring) up to permuting colors. This implies $f\not\sim_{3}g$.
(2) Let $G_{2}$ be the graph as follows:
$G_{2}$
Using Macaulay2, one has $\omega(I_{{\mathcal{M}}_{G_{2}}})=4$. Moreover, for the following two $4$-edge-colorings $f$ and $g$ of $G_{2}$, one has $f\not\sim_{3}g$.
$f$
$g$
(3) Let $G_{3}$ be the graph as follows:
$G_{3}$
Using Macaulay2, one has $\omega(I_{{\mathcal{M}}_{G_{3}}})=4$. Moreover, for the following two $4$-edge-colorings $f$ and $g$ of an edge-replication multigraph of $G_{3}$, one has $f\not\sim_{3}g$. Note that for any two $4$-edge-colorings $f_{1},f_{2}$ of $G_{3}$, it follows that $f_{1}\sim_{4}f_{2}$.
$f$
$g$
(4) Let $G_{4},G_{5},\ldots,G_{8}$ be the graphs as follows:
$G_{4}$
$G_{5}$
$G_{6}$
$G_{7}$
$G_{8}$
Using Macaulay2, one has $\omega(I_{{\mathcal{M}}_{G}})=4$ when $G\in\{G_{4},\ldots,G_{8}\}$. We notice that the graph $G_{7}$ and $G_{8}$ are odd subdivisions of $G_{1}$.
Proposition 2.4 says that, for a graph $G$ and a subgraph $G^{\prime}$ of $G$, one has $\omega(I_{{\mathcal{M}}_{G^{\prime}}})\leq\omega(I_{{\mathcal{M}}_{G}})$ since ${\mathcal{M}}_{G^{\prime}}$ is a face of ${\mathcal{M}}_{G}$. In general, if $G^{\prime}$ is an odd subdivision of a subgraph of $G$, we can obtain the same inequality.
Proposition 6.2.
Let $G$ be a graph, and let $G^{\prime}$ be an odd subdivision of $G$. Then ${\mathcal{M}}_{G^{\prime}}$ has a face that is isomorphic t ${\mathcal{M}}_{G}$. In particular, we have $\omega(I_{{\mathcal{M}}_{G}})\leq\omega(I_{{\mathcal{M}}_{G^{\prime}}})$.
Proof.
Let $E(G)=\{e_{1},\dots,e_{n}\}$ be the edge set of $G$. Suppose that $G^{\prime}$ is obtained from $G$ by replacing an edge $e_{n}$ of $G$ by a path $P=(e_{n}^{\prime},e_{n+1}^{\prime},e_{n+2}^{\prime})$ with three edges. Let $F_{i}$ be a face of ${\mathcal{M}}_{G^{\prime}}$ defined by $F_{i}={\mathcal{M}}_{G^{\prime}}\cap H_{i}$ where
$H_{1}$ $=$ $\{x\in{\mathbb{R}}^{n+2}:x_{n}+x_{n+1}=1\},$
$H_{2}$ $=$ $\{x\in{\mathbb{R}}^{n+2}:x_{n+1}+x_{n+2}=1\}.$
Let $F=F_{1}\cap F_{2}$. Then $F$ is a face of ${\mathcal{M}}_{G^{\prime}}$ and each vertex $x$ of $F$ satisfies $(x_{n},x_{n+1},x_{n+2})\in\{(0,1,0),(1,0,1)\}$. It is easy to see that $(x_{1},\dots,x_{n-1},0,1,0)\in{\mathcal{M}}_{G^{\prime}}$ if and only if $(x_{1},\dots,x_{n-1},0)\in{\mathcal{M}}_{G}$. If $(x_{1},\dots,x_{n-1},1,0,1)\in{\mathcal{M}}_{G^{\prime}}$, then $x_{i}=0$ if $e_{i}$ is adjacent to either $e_{n}^{\prime}$ or $e_{n+2}^{\prime}$ in $G^{\prime}$. Note that $e_{i}$ is adjacent to either $e_{n}^{\prime}$ or $e_{n+2}^{\prime}$ in $G^{\prime}$ if and only if $e_{i}$ is adjacent to $e_{n}$ in $G$. Thus, $(x_{1},\dots,x_{n-1},1,0,1)\in{\mathcal{M}}_{G^{\prime}}$ if and only if $(x_{1},\dots,x_{n-1},1)\in{\mathcal{M}}_{G}$. Let $\varphi:{\mathbb{R}}^{n+2}\rightarrow{\mathbb{R}}^{n+2}$ be a linear transformation defined by $\varphi(x)=(x_{1},\dots,x_{n},x_{n+1}+x_{n},x_{n+2}-x_{n})$. Then $\varphi(F)={\mathcal{M}}_{G}\times\{(1,0)\}\cong{\mathcal{M}}_{G}$.
Since every odd subdivision is obtained from $G$ by replacing an edge of $G$ by a path with three edges repeatedly, we have a desired conclusion from this. ∎
From Propositions 2.1 and 2.4, we have the following.
Corollary 6.3.
Suppose that a graph $G_{0}$ contains an odd subdivision $G^{\prime}$ of a graph $G$ as a subgraph. Then we have $\omega(I_{{\mathcal{M}}_{G}})\leq\omega(I_{{\mathcal{M}}_{G^{\prime}}})\leq\omega(I_{{\mathcal{M}}_{G_{0}}})$.
Hence we obtain the following from Proposition 2.4.
Proposition 6.4.
Let $G$ be a graph and let $G_{1},\ldots,G_{5}$ and $G_{6}$ be the graphs as in Example 6.1. If $G$ contains an odd subdivision of $G_{1},\ldots,G_{5}$ or $G_{6}$ as a subgraph, then one has $\omega(I_{{\mathcal{M}}_{G}})\geq 4$.
Next, we consider Conjecture 1.5. Note that to solve the conjecture, it suffices to consider case of complete graphs from Proposition 2.4. In other words, Conjecture 1.5 is equivalent to the following.
Conjecture 6.5.
Let $K_{d}$ be a complete graph with $d$ vertices. Then one has $\omega(I_{{\mathcal{M}}_{K_{d}}})\leq 4$.
For graphs with a small number of vertices, we consider Conjecture 1.5. Since $K_{4}$ is line perfect and has no $K_{2,3}$ as a subgraph, we have the following from Theorem 5.7.
Corollary 6.6.
Let $G$ be a graph on $[d]$. If $d\leq 4$, then $\omega(I_{{\mathcal{M}}_{G}})=2$.
For graphs with $5$, 66 and $7$ vertices, we verify Conjecture 1.5 through computational experiments. By using Macaulay2, we can confirm that $\omega(I_{{\mathcal{M}}_{K_{7}}})=4$. Hence, for graphs $G$ with $5$, 66 and $7$ vertices, one has $\omega(I_{{\mathcal{M}}_{G}})\leq 4$. Additionally, we classify all graphs $G$ such that $\omega(I_{{\mathcal{M}}_{G}})=4$ by using both Macaulay2 and Nauty [14]. Let $G$ be a graph on $[d]$ and let $G_{1},G_{2},\ldots,G_{8}$ be the graphs as in Example 6.1, respectively. Then one has
  1. (1)
    Assume $d=5$. Then $\omega(I_{{\mathcal{M}}_{G}})=4$ if and only if $G$ contains $G_{1}$ as a subgraph. Otherwise, $\omega(I_{{\mathcal{M}}_{G}})\leq 3$.
  2. (2)
    Assume $d=6$. Then $\omega(I_{{\mathcal{M}}_{G}})=4$ if and only if $G$ contains $G_{1}$ or $G_{2}$ as a subgraph. Otherwise, $\omega(I_{{\mathcal{M}}_{G}})\leq 3$.
  3. (3)
    Assume $d=7$. Then $\omega(I_{{\mathcal{M}}_{G}})=4$ if and only if $G$ contains one of $G_{1},G_{2},\ldots,G_{8}$ as a subgraph. Otherwise, $\omega(I_{{\mathcal{M}}_{G}})\leq 3$.
Next, we consider Conjecture 1.5 for a class of graphs. For $d\geq 4$, let $W_{d}$ be the graph on $[d]$ whose edge set is
$\{\{1,2\},\{2,3\},\ldots,\{d-2,d-1\},\{1,d-1\}\}\cup\{\{1,d\},\{2,d\},\ldots,\{d-1,d\}\}.$
$W_{4}$
$W_{5}$
$W_{6}$
The graph $W_{d}$ is called a wheel graph. Note that $W_{5}$ is the graph $G_{1}$ as in Example 6.1. If $d$ is odd, then $W_{d}$ contains an odd subdivision of $W_{5}$ as a subgraph. In this case, one has $\omega(I_{{\mathcal{M}}_{W_{d}}})\geq 4$. From Corollary 6.6, $\omega(I_{{\mathcal{M}}_{W_{4}}})=2$. Moreover, by using Macaulay2, we obtain $\omega(I_{{\mathcal{M}}_{W_{6}}})=2$. This leads the following conjecture.
Conjecture 6.7.
Let $d$ be an integer $\geq 4$. Then one has
$\omega(I_{{\mathcal{M}}_{W_{d}}})=\begin{cases}2&\mbox{if $d$ is even},\\ 4&\mbox{if $d$ is odd}.\end{cases}$
By computational experiments, we can confirm this conjecture for $d\leq 10$.
Finally, we discuss perfect matching polytopes. Since ${\mathcal{P}}_{G}$ is a face of ${\mathcal{M}}_{G}$ for any graph $G$, we have the following from Theorem 1.3 and Proposition 2.1.
Proposition 6.8.
Let $G$ be a graph with $n$ edges. Then $\omega(I_{{\mathcal{P}}_{G}})\leq r$ if and only if for any ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$ and for any $k$-edge-colorings $f$ and $g$ of $G^{(e)}_{{\mathbf{a}}}$ such that each color of $f$ and $g$ corresponds to a perfect matching of $G$, one has $f\sim_{r}g$.
Proposition 6.9.
Let $G$ be a graph on $[d]$ with $n$ edges. Then we have the following:
  • (a)
    $\omega(I_{{\mathcal{P}}_{G}})\leq\omega(I_{{\mathcal{M}}_{G}})$;
  • (b)
    There exists a graph $G^{\prime}$ on $[2d]$ with $2n+d$ edges such that $\omega(I_{{\mathcal{M}}_{G}})\leq\omega(I_{{\mathcal{P}}_{G^{\prime}}})$.
Proof.
(a) Since ${\mathcal{P}}_{G}$ is a face of ${\mathcal{M}}_{G}$, the assertion follows from Proposition 2.1.
(b) The idea of the proof comes from the argument appearing in e.g., [6, p.129]. Let $G^{\prime}$ be a graph on $[2d]$ obtained by taking two copies of $G$ and adding edges between each vertex and its copy. Let $r=\omega(I_{{\mathcal{P}}_{G^{\prime}}})$ and assume that $r<\omega(I_{{\mathcal{M}}_{G}})$. From Theorem 1.3, there exists ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$ and $k$-edge-colorings $f$ and $g$ of $G^{(e)}_{{\mathbf{a}}}$, such that $f\not\sim_{r}g$. Let ${\mathbf{a}}^{\prime}=({\mathbf{a}},{\mathbf{a}},{\mathbf{b}})\in{\mathbb{Z}}_{\geq 0}^{2n+d}$ where ${\mathbf{b}}=(b_{1},\dots,b_{d})$ with $b_{i}=k-\deg_{G_{\mathbf{a}}}(i)$. Then ${G^{\prime}}^{(e)}_{{\mathbf{a}}^{\prime}}$ is a $k$-regular multigraph. We define a $k$-edge coloring $f^{\prime}$ of ${G^{\prime}}^{(e)}_{{\mathbf{a}}^{\prime}}$ as follows:
  • If $e$ is an edge of one of copies of $G$, then we set $f^{\prime}(e)=f(e)$.
  • If $e_{1},\dots,e_{s}$ with $s=k-\deg_{G_{\mathbf{a}}}(i)$ are edges between the vertex $i$ and its copy, then colors $f^{\prime}(e_{1}),\dots,f^{\prime}(e_{s})$ are distinct each other and different from $\{f(e):e\mbox{ is adjacent to }i\mbox{ in }G\}$.
Similarly, we define $g^{\prime}$ from $g$. Then each color of $f^{\prime}$ (resp. $g^{\prime}$) is a perfect matching of $G^{\prime}$. From Proposition 6.8, we have $f^{\prime}\sim_{r}g^{\prime}$. Since $G^{(e)}_{\mathbf{a}}$ is an induced subgraph of ${G^{\prime}}^{(e)}_{{\mathbf{a}}^{\prime}}$, it follows that $f\sim_{r}g$, a contradiction. ∎
Thus we can obtain the following.
Proposition 6.10.
Conjecture 1.5 is equivalent to Conjecture 1.6.
Appendix A Flow polytopes
Let $Q$ be a directed graph with the vertex set $Q_{0}$ and the arrow set $Q_{1}$. For an arrow $a\in Q_{1}$, let $a^{-}$ be the starting vertex of $a$, and let $a^{+}$ be the terminating vertex of $a$. Given an integer vector $\theta\in{\mathbb{Z}}^{Q_{0}}$ and non-negative integer vectors ${\bf l},{\bf u}\in{\mathbb{Z}}_{\geq 0}^{Q_{1}}$, the flow polytope associated with $Q,\theta,{\bf l},{\bf u}$ is the polytope
$\nabla(Q,\theta,{\bf l},{\bf u})=\left\{x\in{\mathbb{R}}^{Q_{1}}:{\bf l}\leq x\leq{\bf u},\theta(v)=\sum_{a^{+}=v}x(a)-\sum_{a^{-}=v}x(a)\mbox{ for }\forall v\in Q_{0}\right\}.$
We see that the matching polytope of a bipartite graph is a flow polytope.
Proposition A.1.
For any bipartite graph $G$, the matching polytope ${\mathcal{M}}_{G}$ is isomorphic to a flow polytope.
Proof.
Let $G$ be a bipartite graph on the vertex set $V=V_{1}\sqcup V_{2}$ and the edge set $E$. It is known that the matching polytope $\mathcal{M}_{G}$ of $G$ coincides with
$\left\{x\in{\mathbb{R}}^{E}:\begin{array}[]{cl}x(e)\geq 0&\mbox{for }\forall e\in E\\ \sum_{e\ni v}x(e)\leq 1&\mbox{for }\forall v\in V\end{array}\right\}.$
Hence we have
${\mathcal{M}}_{G}=\left\{x\in{\mathbb{R}}^{E}:\begin{array}[]{cl}x(e)\geq 0&\mbox{for }\forall e\in E\\ y(v)+\sum_{e\ni v}x(e)=1&\mbox{for }\forall v\in V\\ y(v)\geq 0&\mbox{for }\forall v\in V\end{array}\right\}.$
Thus $\mathcal{M}_{G}$ is a projection of the polytope
$\mathcal{P}=\left\{(x,y)\in{\mathbb{R}}^{E\cup V}:\begin{array}[]{cl}x(e)\geq 0&\mbox{for }\forall e\in E\\ y(v)+\sum_{e\ni v}x(e)=1&\mbox{for }\forall v\in V\\ y(v)\geq 0&\mbox{for }\forall v\in V\end{array}\right\}.$
Note that, if $(x,y)\in{\mathcal{P}}$, then we have
$\sum_{v\in V_{2}}y(v)-\sum_{v\in V_{1}}y(v)=\sum_{v\in V_{2}}\left(1-\sum_{e\ni v}x(e)\right)-\sum_{v\in V_{1}}\left(1-\sum_{e\ni v}x(e)\right)=|V_{2}|-|V_{1}|.$
Hence $\mathcal{P}$} is the flow polytope $\nabla(Q,\theta,{\bf 0},{\bf 1})$ where $Q$ is the directed graph on the vertex set $Q_{0}=V_{1}\cup V_{2}\cup\{v_{0}\}$ and the arrow set
$Q_{1}=\{(i,j):i\in V_{1},j\in V_{2}\}\cup\{(i,v_{0}):i\in V_{1}\}\cup\{(v_{0},j):j\in V_{2}\},$
and $\theta=(-1,\dots,-1,1,\dots,1,|V_{2}|-|V_{1}|)\in{\mathbb{Z}}^{V_{1}\cup V_{2}\cup\{v_{0}\}}$. Since ${\mathcal{M}}_{G}$ is a projection of $\mathcal{P}$} with
$\dim{\mathcal{P}}=|Q_{1}|-|Q_{0}|+1=(|E|+|V|)-(|V|+1)+1=|E|=\dim{\mathcal{M}}_{G},$
it follows that ${\mathcal{M}}_{G}$ is isomorphic to ${\mathcal{P}}=\nabla(Q,\theta,{\bf 0},{\bf 1})$. ∎
Acknowledgment
This work was supported by Grant-in-Aid for JSPS Fellows 23KJ2117, JSPS KAKENHI 24K00534 and 22K13890.